OpenStudy (anonymous):

how do you get answer for this question Rhonda has 1.35 in nickels and dimes in her pocket. if she has six more dimes than nickels,which equation can be used to determine x, the number of nickels she have?

OpenStudy (anonymous):

\(\Huge \sf \hspace{10 pt}{\color{black}{Welcome~to~OpenStudy!!}} \hspace{-259 pt} \color{#ff7100}{Welcome}\ \color{#999900}{to} \hspace{6 pt} \color{#00B4ff}{Open}\color{#7cc517}{Study!!}\\ \hspace{100 pt}\sf \ by\ DataLg2 \) Since you're a new user, you can watch the \(\sf \href{http:///www.youtube.com/watch?v=lDiiH8PItto}{OS\ tutorial}\) made by wolfe8. As a new user, we expect that you already read the \(\sf \href{http:///openstudy.com/terms-and-conditions}{Terms\ and\ Conditions}\) or TaC before signing up. Here in Openstudy, we keep the good and friendly atmosphere by following TaC and \(\sf \href{http:///openstudy.com/code-of-conduct}{OS's\ Code\ of\ Conduct}\) (CoC), which I also recommend you to read in order to avoid having troubles with this site. If you have more questions about openstudy, feel free to ask us at the \(\sf \href{http:///openstudy.com/study#/groups/Openstudy%20Feedback}{OpenStudy\ Feedback}\). Once again, welcome and enjoy studying here at \(\sf \color{#00B4ff}{Open}\color{#7cc517}{Study!!}\)! Now with your problem, let x be the number of nickels y be the number of dimes From this statement: "Rhonda has 1.35 in nickels and dimes in her pocket." we know that we will have \(\sf x + y = 1.35\) as our first equation. "he has six more dimes than nickels" which means we will have \(\sf y = x + 6\) as our second equation so that's what we can do for now since your question seems incomplete. If you could kindly take a screenshot of your questin or type the complete qeusion, we can figure it out the anwer together (:

OpenStudy (anonymous):

*question , sorry for the typo ;p

OpenStudy (anonymous):

The next thing we can do is to combine those two equations together to solve for x by either substitution method or elimination method.

OpenStudy (anonymous):

okay, so let's use substitution method. @touchone I need your participation in evaluating this ok? so we have: \(\sf x + \color{red}{y} = 1.35\) -> eqn 1 \(\sf \color{red}{y = x + 6}\) -> eqn 2 replace \(\sf \color{red}{y}\) from eqn 1, using eqn can you do that for me?

OpenStudy (anonymous):

**eqn 1 should be 0.5x + 0.10 y =1.35 not x + y = 1.35 sorry my bad

OpenStudy (anonymous):